1. Kylie Jenner, a beautiful model & actress, is also a businesswoman. She sells two commodities (product x and product y) that are a hit in the market. There is only one problem; she could not figure out if her products are substitutes or if they complement each other. Commodity X sold around 1000 pieces originally. Commodity X had a new sales volume of 1200 the next month. The price of X was constant. Commodity Y had an original price of P50 but was later increased to P100. Y originally sold around 900 pieces last month but since its price increased, its demand went down to 700. Airplane tickets increased by 10% and movie tickets increased around 50%. Kylie was depressed and wanted to stop being an actress to be a full-time businesswoman. Compute for cross elasticity of demand. Are the commodities substitutes or complementary? cross elasticity of demand coefficient.
1
Expert's answer
2016-07-24T11:52:02-0400
QX1 = 1000 pieces, QX2 = 1200, PX was constant. PY1 = 50, PY2 = 100, QY1 = 900 pieces, QY2 = 700.
The cross elasticity of demand is:
Ed = (1200 - 1000)/(100 - 50)*(100 + 50)/(1200 + 1000) = 200/50*150/2200 = 6/22.
As Ed > 0, so the commodities are substitutes.